OpenStudy (anonymous):

A rectangle has a length of 15 cm and a perimeter of 72 cm. Which equation could be used to find the width of the rectangle? A. 2w + 15 = 72 B. w + 15 = 72 C. 2w + 30 = 72 D. 15w = 72

OpenStudy (igreen):

The formula for the Perimeter of a rectangle is P = 2L + 2W. Plug in what we know: Perimeter = 72 Length = 15 P = 2L + 2W 72 = 2(15) + 2W

OpenStudy (igreen):

You can simplify it more: 72 = 2(15) + 2W 72 = 30 + 2W

OpenStudy (igreen):

Therefore, the answer is C.

OpenStudy (anonymous):

yes, what @iGreen said. Here maybe this video can help a little bit: https://www.khanacademy.org/math/geometry/basic-geometry/perimeter_area_tutorial/v/perimeter-and-area-basics

OpenStudy (anonymous):

ty\

OpenStudy (anonymous):

ok wat time is it over there kitkat

OpenStudy (anonymous):

i got a 100%

OpenStudy (anonymous):

@Devil it is 10:14 am, I can't webcall you right now but I can at 1:50 p.m. to 2:00 p.m. my time for like 15 minutes

OpenStudy (anonymous):

mine is 11:16

OpenStudy (anonymous):

@Devilwolf3000 ok, so you are an hour ahead, how bout I pm you when I am ready to webcall you and you can send me the link to the tinychat thingy

OpenStudy (anonymous):

what a minute were you the one that i saw when i tinychat teddy like a month ago mine name was austin_marshall

OpenStudy (anonymous):

@Devilwolf3000 no I just found OS on Friday....

OpenStudy (anonymous):

ohhh kk fan me so you can pm me kk

OpenStudy (anonymous):

@Devilwolf3000 ok, but when I get on again my username is gonna b a little Different kk

OpenStudy (anonymous):

what will it be when i see you

OpenStudy (anonymous):

Idk, probably something along the lines of this user name just a little different I will tag you kk?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

You can close this question now if u want ;)

OpenStudy (igreen):

Medal? c:

OpenStudy (anonymous):

sorry i was doin social studies forgot to close but i got an 100%

OpenStudy (anonymous):

there